### v2

# Spontaneous Decoherence from Imaginary–Order Spectral Deformations

Abstract
A mechanism of spontaneous decoherence is examined in which the generator of quantum dynamics is replaced by the imaginary–order - fundamentally different from real-order fractional calculus - spectral deformation $H^{1+i\beta}$ for a positive self–adjoint Hamiltonian $H$. The deformation modifies dynamical phases through the factor $E^{i\beta}=e^{i\beta\log E}$, whose rapid oscillation suppresses interference between distinct energies. A non–stationary–phase analysis yields quantitative estimates: oscillatory contributions to amplitudes or decoherence functionals decay at least as $\mathcal{O}(1/|\beta|)$. The kinematical structure of quantum mechanics—the Hilbert–space inner product, projection operators, and the Born rule—remains unchanged; the modification is entirely dynamical and acts only through spectral phases.
Physical motivations for the deformation arise from clock imperfections, renormalization–group and effective–action corrections that introduce logarithmic spectral terms, and semiclassical gravity analyses in which complex actions produce spectral factors of the form $E^{i\beta}$.
The mechanism is illustrated in examples relevant to quantum–gravity–inspired quantum mechanics, including two–level systems, quartic oscillators, FRW minisuperspace toy models, and effective one–dimensional curved–background Hamiltonians. In each case the spectral structure yields explicit decoherence rates under the deformation. The parameter $\beta$ may, in principle, be experimentally constrained through precision coherence measurements in low–noise quantum platforms. As a benchmark, we estimate bounds $|\beta|\lesssim 10^{-5}$ from current superconducting–qubit coherence times, and discuss how other platforms such as trapped ions, NV centers, and cold atoms, with longer coherence times, can further strengthen sensitivity.
A detailed related–work analysis contrasts the present mechanism with Milburn–type intrinsic decoherence, Diósi–Penrose gravitational collapse, GRW/CSL models, clock–induced decoherence, and energy–conserving collapse models, as well as environmental frameworks such as Lindblad master equations, Caldeira–Leggett baths, and non–Hermitian Hamiltonian deformations. None of these approaches produces decoherence through a purely unitary imaginary–order spectral deformation of a single positive Hamiltonian. This positions $H^{1+i\beta}$ dynamics as a compact, testable, and genuinely novel phenomenological encapsulation of logarithmic spectral corrections arising in quantum–gravity– motivated effective theories, while remaining fully compatible with standard quantum kinematics.
1 Introduction
Let $H$ be a positive self–adjoint Hamiltonian acting on a Hilbert space. Standard quantum dynamics is generated by the unitary one–parameter group
$U(t)=e^{-iHt},$ (1)
with interference arising from coherent superpositions of distinct spectral components. In conventional treatments, decoherence is usually attributed to environmental coupling, coarse–graining, or stochastic modifications of Schrödinger dynamics. Explicit models include Lindblad master equations and Caldeira–Leggett oscillator baths for open systems [8, 9], as well as intrinsic or spontaneous decoherence proposals that modify the dynamics directly, such as Milburn’s stochastic time–step model [1], Diósi–Penrose gravitational collapse [2, 3], GRW/CSL wave–function reduction [4, 5, 6], and clock–induced decoherence in quantum gravity–motivated settings [7]. More recently, energy–conserving spontaneous collapse models have been explored by Snoke and collaborators [12, 13, 14, 15]. In parallel, non–Hermitian Hamiltonian deformations have been used to model open–system effects and time–keeping imperfections in a controlled way [10, 11].
Point of departure.
This paper investigates an entirely different mechanism in which decoherence arises spontaneously from a deterministic modification of the dynamical generator rather than from external degrees of freedom or stochastic processes. The proposal is to replace $H$ by the imaginary–order spectral deformation $H^{1+i\beta}$, defined through the spectral theorem. The evolution operator becomes
[…]
with eigenstate amplitudes $e^{-itE^{1+i\beta}}$. The exponent introduces phases depending on $\beta\log E$, and the resulting oscillatory structure suppresses interference between distinct energies. The mechanism is deterministic and internal to the system: it does not rely on baths, measurements, or nonlinearity. The Hilbert–space inner product, projection operators, and probability assignments via the Born rule remain exactly as in standard quantum mechanics. Only the dynamical phases are modified.
A poetic intuition.
Echoing Wheeler’s famous dictum that space tells matter how to move and matter tells space how to curve, it is tempting to summarize the spirit of the mechanism in a single line:
Energy tells time how to tick; time tells energy how to decohere.
This line is intended as an intuition rather than as a literal dynamical feedback loop. In the formalism, time remains the usual external evolution parameter, exactly as in ordinary quantum mechanics. What actually “carries” the effect is the evolution operator
[…]
under which each energy eigencomponent acquires an additional phase
$E^{i\beta}=e^{i\beta\log E}.$ (4)
Decoherence arises because these phases diverge across the spectrum, leading to suppression of interference by non–stationary–phase cancellation rather than by dissipation or stochastic noise. The imagery captures the qualitative idea that different energies “keep time” in slightly different ways, while the underlying mechanism is precise: the spectral components of the state accumulate energy–dependent phases, and it is their mutual dephasing that suppresses coherence.
Scope of this work.
The present proposal is intended as a phenomenological mechanism for spontaneous decoherence within strictly unitary quantum dynamics, encoded by the imaginary–order spectral deformation $H^{1+i\beta}$. It does not aim to solve the measurement problem, to replace standard environment–induced decoherence, or to derive $\beta$ from a specific microscopic or quantum–gravity model. Rather, it provides a compact way to parameterize logarithmic spectral corrections and to assess their potential impact on coherence in realistic platforms.
Structure of paper.
Section 2 introduces imaginary–order spectral deformations and establishes basic properties. Section 3 presents a non–stationary–phase estimate that quantifies interference suppression. Section 4 develops physical motivations from clock imperfections, renormalization–group flow, semiclassical gravity, and curved–spacetime field theory. Section 5 explains how the present mechanism differs from standard real–order fractional calculus and fractional quantum mechanics. In Section 6, we illustrate the mechanism in several simple Hamiltonians. Section 7 discusses phenomenology and experimental constraints, including a numerical estimate for superconducting qubits and a summary table of multiple experimental platforms. Section 8 situates the proposal among intrinsic and environmental decoherence models and highlights its novelty. Section 9 presents a brief discussion and outlook.
2 Imaginary–Order Spectral Deformations
Let $H$ be a positive self–adjoint operator with spectral resolution
$H=\int_{0}^{\infty}E\,d\Pi(E),$ (5)
where $\Pi(E)$ is the projection–valued measure associated with $H$. For any complex exponent $\alpha$, the functional calculus defines
$H^{\alpha}=\int_{0}^{\infty}E^{\alpha}\,d\Pi(E),$ (6)
where $E^{\alpha}$ is defined using the principal branch of the logarithm: for $E>0$ and $\alpha=1+i\beta$ with $\beta\in\mathbb{R}$,
$E^{1+i\beta}=E\,e^{i\beta\log E}.$ (7)
In this paper we focus on the deformation
$H\mapsto H^{1+i\beta}=\int_{0}^{\infty}E^{1+i\beta}\,d\Pi(E),$ (8)
and the corresponding evolution operator
$U_{\beta}(t)=\exp(-itH^{1+i\beta})=\int_{0}^{\infty}e^{-itE^{1+i\beta}}\,d\Pi(E).$ (9)
Since $H$ is positive and self–adjoint, $H^{1+i\beta}$ is a closed normal operator with domain $D(H^{1+i\beta})=D(H)$, and the spectral decomposition implies
[…]
The norm is preserved:
$U_{\beta}(t)^{\dagger}U_{\beta}(t)=\int_{0}^{\infty}e^{+itE^{1-i\beta}}\,d\Pi(E)\int_{0}^{\infty}e^{-itE^{1+i\beta}}\,d\Pi(E)=\mathbbm{1},$ (11)
so the dynamics is unitary even though decoherence occurs in the sense of suppression of interference terms.
3 Non–Stationary Phase and Decoherence
Consider an amplitude of the form
$I_{\beta}(t)=\int_{0}^{\infty}f(E)\,e^{-itE^{1+i\beta}}\,dE,$ (12)
where $f$ is a smooth function with compact support in $(0,\infty)$. Define the phase
$\Phi_{\beta}(E)=tE^{1+i\beta},\qquad\Phi_{\beta}^{\prime}(E)=t(1+i\beta)E^{i\beta}.$ (13)
Since $|E^{i\beta}|=1$, we have
$|\Phi_{\beta}^{\prime}(E)|=|t|\,\sqrt{1+\beta^{2}}\geq c|t||\beta|\quad\text{for }|\beta|\geq 1$ (14)
with some constant $c>0$ independent of $E$.
Theorem 1.
Let $f\in C_{c}^{1}((0,\infty))$ and $t\neq 0$. Then there exist constants $C,\beta_{0}>0$ such that for all $|\beta|\geq\beta_{0}$,
$|I_{\beta}(t)|\leq\frac{C}{|\beta|}.$ (15)
In particular, $I_{\beta}(t)\to 0$ as $|\beta|\to\infty$ with decay at least of order $1/|\beta|$.
Proof.
Assume $\operatorname{supp}f\subset[E_{\min},E_{\max}]\subset(0,\infty)$. Integrating by parts,
$I_{\beta}(t)=\left[\frac{f(E)}{-i\Phi_{\beta}^{\prime}(E)}e^{-i\Phi_{\beta}(E)}\right]_{E_{\min}}^{E_{\max}}+\int_{E_{\min}}^{E_{\max}}\frac{f^{\prime}(E)}{i\Phi_{\beta}^{\prime}(E)}e^{-i\Phi_{\beta}(E)}\,dE.$ (16)
Using $|\Phi_{\beta}^{\prime}(E)|\geq c|t||\beta|$ for $|\beta|\geq 1$ and the boundedness of $f$ and $f^{\prime}$, we obtain
$|I_{\beta}(t)|\leq\frac{\|f\|_{\infty}}{c|t||\beta|}+\frac{\|f^{\prime}\|_{\infty}(E_{\max}-E_{\min})}{c|t||\beta|}\leq\frac{C}{|\beta|}$ (17)
for a suitable constant $C>0$ independent of $\beta$. ∎
The theorem shows that oscillatory integrals involving $E^{1+i\beta}$ phases are suppressed as $|\beta|\to\infty$. For interference between two discrete energy levels $(E_{m},E_{n})$, a natural effective decoherence timescale is
$\tau_{\mathrm{dec}}(E_{m},E_{n};\beta)\sim\frac{1}{|\beta|\,\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|},$ (18)
so that interference between those levels is strongly suppressed once $|t|\gg\tau_{\mathrm{dec}}$.
4 Physical Motivation
In this section we sketch several contexts in which a deformation effectively of the form $H\mapsto H^{1+i\beta}$ arises or is suggested.
4.1 Clock imperfections and operational time
A physical clock is described by its own Hamiltonian $H_{C}$. If its tick rate depends slightly on energy, conditioning the system on the clock can introduce phases with logarithmic corrections in the effective system evolution. Schematically, one may obtain factors of the form
$\exp[-it\,E\big]\;\longrightarrow\;\exp[-it\,E\,f(\log E)\big],$ (19)
with $f(\log E)\approx 1+i\beta$ over some energy window. This leads directly to an effective generator $H^{1+i\beta}$ in the operational description of time evolution. In such a setting, $\beta$ encodes small deviations of physical time from the ideal external parameter appearing in Schrödinger evolution. Related ideas appear in clock–induced decoherence and fundamental limits to timekeeping [7].
4.2 Renormalization group and effective actions
Renormalization–group and effective–action treatments of quantum field theories and quantum gravity frequently produce running couplings and masses of the form
$g(E)=g_{0}+\alpha\log(E/\mu),$ (20)
or effective Hamiltonians
$H_{\mathrm{eff}}(E)=Ef(\log E),$ (21)
with $f$ a slowly varying function. When these expressions enter the time–evolution factor $\exp[-itH_{\mathrm{eff}}(E)]$, the logarithmic dependence generates spectral factors of the type $E^{i\beta}$. The deformation $H^{1+i\beta}$ can thus be viewed as a compact representation of RG–induced distortions of the generator, capturing the leading logarithmic behavior in a simple analytic form.
4.3 Semiclassical gravity and complex actions
Anomaly–induced effective actions and semiclassical gravitational calculations often involve operators such as $\log(\Box/\mu^{2})$, where $\Box$ is a covariant d’Alembertian. In minisuperspace or reduced models, these terms translate into logarithmic functions of the effective Hamiltonian or energy. Standard references on quantum fields in curved spacetime [18, 19] contain numerous examples in which logarithmic spectral terms appear.
Likewise, semiclassical WKB analyses of gravitational systems and black–hole spacetimes can yield actions of the form
$S(E)=S_{0}(E)+i\hbar\,\gamma\log\left(\frac{E}{E_{P}}\right)+\cdots,$ (22)
so that the semiclassical wave function
$\exp\left(\frac{i}{\hbar}S(E)\right)=E^{i\gamma}\,\exp\left(\frac{i}{\hbar}S_{0}(E)\right)$ (23)
contains precisely the spectral factor responsible for the deformation studied here.
In this perspective, $H^{1+i\beta}$ should be interpreted as a phenomenological stand–in for a more general functional deformation $F(H)$ arising from integrating out gravitational or high–energy degrees of freedom, with the particular power $1+i\beta$ capturing the leading logarithmic behavior without committing to a specific microscopic model.
5 Imaginary-order (not Real-order Fractional) Calculus
The appearance of a complex exponent $1+i\beta$ might invite comparison with fractional calculus and fractional quantum mechanics [16, 17]. In standard fractional calculus, one studies real–order derivatives and integrals (e.g. Riemann–Liouville or Caputo derivatives) that act via nonlocal integral kernels and encode memory effects. Real–order fractional dynamics often lead to anomalous diffusion and non–Markovian behavior.
The present construction is of a different type. The deformation
$H^{1+i\beta}=\int_{0}^{\infty}E^{1+i\beta}\,d\Pi(E)$ (24)
belongs to the spectral functional calculus of a positive operator $H$. There is no associated real–order memory kernel in time and no direct connection to the Riemann–Liouville or Caputo operators. The fractional exponent appears in the energy domain, not as a fractional time derivative. The resulting dynamics remain unitary, and decoherence arises from oscillatory phase suppression rather than from anomalous diffusion.
In particular, the mechanism here should not be viewed as an instance of fractional Schrödinger dynamics. It is instead a purely spectral deformation in the sense of functional calculus, whose primary consequence is a dephasing of energy eigencomponents governed by $E^{i\beta}$.
6 Illustrative Examples
We now illustrate how the deformation $H^{1+i\beta}$ acts on several simple Hamiltonians. The goal is not exhaustive analysis, but to show concretely how spectral structure feeds into decoherence rates.
6.1 Two–level system
Consider a two–level system with Hamiltonian
$H=\begin{pmatrix}E_{1}&0\\ 0&E_{2}\end{pmatrix},\qquad 0<E_{1}<E_{2}.$ (25)
The deformed generator is
$H^{1+i\beta}=\begin{pmatrix}E_{1}^{1+i\beta}&0\\ 0&E_{2}^{1+i\beta}\end{pmatrix}$, (26)
and the evolution operator
$U_{\beta}(t)=\begin{pmatrix}e^{-itE_{1}^{1+i\beta}}&0\\ 0&e^{-itE_{2}^{1+i\beta}}\end{pmatrix}$. (27)
[…]
$\rho_{12}(t)=c_{1}\overline{c_{2}}\,e^{-it\left(E_{1}^{1+i\beta}-E_{2}^{1+i\beta}\right)}.$ (28)
The phase difference can be written as
$E_{1}^{1+i\beta}-E_{2}^{1+i\beta}=E_{1}e^{i\beta\log E_{1}}-E_{2}e^{i\beta\log E_{2}},$ (29)
and the non–stationary–phase analysis implies that superpositions involving many such modes exhibit $1/|\beta|$ suppression of interference. Even in the two–level case, the effective decoherence timescale can be estimated from $\tau_{\mathrm{dec}}\sim 1/(|\beta|\,|E_{1}\log E_{1}-E_{2}\log E_{2}|)$.
6.2 Quartic oscillator
As a simple bound–state example, consider the quartic oscillator
$H=\frac{p^{2}}{2m}+\lambda x^{4},\qquad\lambda>0.$ (30)
Its eigenvalues satisfy asymptotically
$E_{n}\sim\kappa\,n^{4/3},\qquad n\to\infty,$ (31)
[…]
$\tau_{\mathrm{dec}}(n,m;\beta)\sim\frac{1}{|\beta|\,\big|E_{n}\log E_{n}-E_{m}\log E_{m}\big|}.$ (32)
For large $n,m$, the asymptotic form of $E_{n}$ can be used to obtain explicit scaling relations between decoherence time, level index, and $\beta$.
6.3 FRW minisuperspace toy model
In FRW minisuperspace models with a single scale factor $a$, effective Hamiltonians of the form
$H=-\frac{d^{2}}{da^{2}}+V(a)$ (33)
arise, with potentials $V(a)$ encoding curvature, cosmological constant, and matter content. For simple choices of $V(a)$ (e.g. harmonic–oscillator–like near a minimum), one obtains a discrete spectrum $E_{n}$ that grows approximately linearly in $n$. The deformation $H^{1+i\beta}$ then introduces phases $E_{n}^{1+i\beta}$ that dephase superpositions of minisuperspace modes, providing a toy model of spontaneous decoherence in a cosmological setting.
6.4 Curved–background and Schwarzschild interior–type Hamiltonians
Effective one–dimensional Hamiltonians modeling motion in curved backgrounds, including toy models of the Schwarzschild interior, take the form
$H=-\frac{d^{2}}{dx^{2}}+V(x),$ (34)
with potentials such as
$V(x)=-\frac{\alpha}{x^{2}}+\beta_{1}x^{2},\qquad\alpha,\beta_{1}>0.$ (35)
This class of potentials admits a discrete spectrum for appropriate boundary conditions and has been used as a simple model of black–hole interior dynamics. Let $E_{n}$ denote the eigenvalues. Under the deformation, the time evolution of a superposition of interior modes is governed by
[…]
and the interference structure can be studied by approximating the sum as an integral over a smooth spectral density. The absence of stationary points in the deformed phase implies $\mathcal{O}(1/|\beta|)$ suppression of cross terms, providing a simple model of spontaneous decoherence in an effective black–hole interior Hamiltonian.
7 Phenomenology and Experimental Constraints
7.1 General scaling and decoherence envelopes
For a pair of energy levels $(E_{m},E_{n})$ the effective decoherence rate suggested by the non–stationary–phase analysis is
$\Gamma_{mn}(\beta)\sim|\beta|\,\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|,$ (37)
with corresponding timescale $\tau_{\mathrm{dec}}\sim 1/\Gamma_{mn}$. If environmental decoherence produces a standard coherence envelope $C_{\mathrm{std}}(t)$, the present mechanism can be viewed as supplying an additional deterministic decay envelope, so that
$C_{\mathrm{meas}}(t)\approx C_{\mathrm{std}}(t)\,e^{-\Gamma_{mn}t},$ (38)
at least over timescales where the non–stationary–phase estimate is valid. In an experiment, one would first calibrate $C_{\mathrm{std}}(t)$ using environmental models and then fit the residual decay to extract or bound $\Gamma_{mn}$, and hence $\beta$. This logic parallels the way collapse–model parameters are bounded by precision experiments [6].
7.2 Superconducting qubits as a benchmark platform
Consider a superconducting–qubit platform where typical transition energies lie in the range $E\sim 5$ GHz and coherence times $T_{2}\sim 100~\mu\text{s}$ are routinely achieved. In what follows, we measure energies in angular–frequency units (setting $\hbar=1$), so that $E$ and $\Delta E$ are understood as angular frequencies in s-1; numerically, $E\sim 5~\text{GHz}$ should thus be read as $E\sim 2\pi\times 5\times 10^{9}~\text{s}^{-1}$. For two energy levels separated by $\Delta E\sim 100$ MHz, the quantity
$\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|$ (39)
can be approximated by
$\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|\approx|\Delta E|\,|\log E|,$ (40)
with $\log E\approx\log(5\times 10^{9})\approx 22.3$ (in units where $\hbar=1$ and energies are expressed as angular frequencies). Using $\Delta E=10^{8}\,\text{s}^{-1}$, we obtain
$\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|\approx 2.2\times 10^{9}\,\text{s}^{-1}.$ (41)
Requiring spontaneous decoherence from the deformation to occur on timescales longer than the experimentally observed coherence time, i.e. $\tau_{\mathrm{dec}}\gtrsim T_{2}$, gives
$\frac{1}{|\beta|\,|E_{m}\log E_{m}-E_{n}\log E_{n}|}\gtrsim T_{2},$ (42)
and hence
$|\beta|\lesssim\frac{1}{T_{2}\,|E_{m}\log E_{m}-E_{n}\log E_{n}|}\sim\frac{1}{(10^{-4}\,\text{s})(2.2\times 10^{9}\,\text{s}^{-1})}\approx 5\times 10^{-6}.$ (43)
Thus, for superconducting qubits, the absence of anomalous decoherence at the $T_{2}\sim 100~\mu\text{s}$ level would constrain the deformation parameter to
$|\beta|\lesssim 10^{-5}.$ (44)
Platforms with longer coherence times, such as trapped ions or certain cold–atom setups, can in principle provide even stronger bounds.
Operationally, a practical way for an experimentalist to fit $\beta$ is to treat the deformation as supplying an additional deterministic decay envelope on top of the usual environmental coherence function. If $C_{\mathrm{std}}(t)$ denotes the standard coherence curve obtained from Ramsey or spin–echo measurements, the prediction of the present mechanism is that the measured signal can be modeled as
$C_{\mathrm{meas}}(t)=C_{\mathrm{std}}(t)\,e^{-\Gamma_{mn}t},\qquad\Gamma_{mn}=|\beta|\,\big|E_{n}\log E_{n}-E_{m}\log E_{m}\big|.$ (45)
In practice, one fits the residual exponential envelope after independently calibrating $C_{\mathrm{std}}(t)$; the slope of the residual decay yields $\Gamma_{mn}$, and thus $\beta$, directly. If no statistically significant residual decay is observed, the sensitivity of the fit provides an upper bound on $|\beta|$. In this way, routine coherence measurements become quantitative probes of logarithmic spectral deformations.
7.3 Summary table of representative platforms
It is useful to summarize, at a qualitative level, how different experimental platforms map into sensitivity to the deformation parameter $\beta$. For each, one may identify representative ranges of transition frequencies $E$, level splittings $\Delta E$, and coherence times $T_{2}$, and then apply the scaling
$|\beta|_{\mathrm{max}}\sim\frac{1}{T_{2}\,|\Delta E|\,|\log E|},$ (46)
obtained by approximating $\big|E_{m}\log E_{m}-E_{n}\log E_{n}\big|\approx|\Delta E|\,|\log E|$.
Table 1 lists illustrative parameter regimes and the corresponding order–of–magnitude sensitivity to $|\beta|$. The numbers are indicative and not tied to any particular experiment; they are meant to convey the relative leverage of different platforms given the same underlying scaling (46).
Platform Typical $E$ (Hz) $\Delta E$ (Hz) $T_{2}$ (s) Illustrative $|\beta|_{\mathrm{max}}$
Superconducting qubits $\sim 5\times 10^{9}$ $\sim 10^{8}$ $\sim 10^{-4}$ $\sim 10^{-5}$
Trapped ions (optical qubits) $\sim 10^{15}$ $\sim 10^{6}$$0^{7} 1\sim 1$ $\sim 10^{-7}$$10^{-8}$
NV centers / solid–state spins $\sim 3\times 10^{9}$ $\sim 10^{7}$$10^{8}$ $\sim 10^{-3}$$10^{-2}$ $\sim 10^{-6}$$10^{-7}$
Cold atoms in optical lattices $\sim 10^{4}$$10^{5}$ $\sim 10^{2}$$0^{3} 1\sim 1$1010 $\sim 10^{-5}$$10^{-6}$
Table 1: Representative order–of–magnitude parameters for several experimental platforms and corresponding illustrative sensitivities to the deformation parameter $|\beta|$, based on the scaling (46). The values are indicative and are not tied to any specific laboratory implementation.
The table highlights two qualitative trends. First, longer coherence times $T_{2}$ directly strengthen sensitivity to $|\beta|$. Second, larger transition energies $E$ increase $|\log E|$, and moderate splittings $\Delta E$ in such regimes can further enhance the factor $|\Delta E\,\log E|$. Platforms combining large $T_{2}$ and sizable $|\Delta E\log E|$ are thus especially promising for constraining or detecting the effects of imaginary–order spectral deformations.
8 Related Work and Novelty
This section situates the present proposal among existing models of decoherence, both non–environmental (intrinsic) and environmental (open–system) in character.
8.1 Non–environmental (intrinsic) models
Several proposals have attempted to describe intrinsic or spontaneous decoherence without explicit environments.
Milburn’s intrinsic decoherence model replaces smooth Schrödinger evolution by a stochastic sequence of unitary kicks characterized by a mean waiting time, leading to a master equation of Lindblad form [1]. Diósi and Penrose argue for gravitational collapse of the wave function based on gravitational self–energy considerations, yielding a characteristic collapse timescale of order $\hbar/E_{\Delta}$ [2, 3]. GRW and CSL models implement spontaneous localization events driven by noise, leading to fundamentally nonunitary dynamics [4, 5, 6]. Clock–induced decoherence in quantum gravity–motivated settings attributes decoherence to the use of physical (imperfect) clocks in place of an ideal external time parameter, leading again to Lindblad–type terms and exponential damping of coherences [7].
More recently, Snoke has developed energy–conserving spontaneous collapse models [12], extended to nonlocal relativistic quantum field theory [13] and elaborated with explicit experimental predictions [14] and a broader interpretive framework [15]. In these models, collapse occurs in an energy–conserving manner, but the dynamics remains stochastic and nonunitary at the level of state evolution.
All of these intrinsic models introduce nonunitarity through stochastic processes, collapse operators, or explicit Lindblad terms. None arise from a unitary spectral deformation of the generator in the sense of $H\mapsto H^{1+i\beta}$.
Model Mechanism Comments
Milburn intrinsic decoherence [1] Discrete time steps; stochastic unitary kicks; Lindblad–type master equation Nonunitary; exponential decoherence; no spectral functional calculus.
Diósi–Penrose collapse [2, 3] Gravitational self–energy induces collapse Stochastic; modifies state norm; not phase–only.
GRW/CSL [4, 5, 6] Spontaneous localization events Noise–driven; fundamentally nonunitary.
Clock–induced decoherence [7] Time uncertainty induces additional Lindblad terms Exponential damping; no imaginary–order spectral power.
Snoke energy–conserving collapse [12, 13, 14, 15] Spontaneous collapse with energy conservation Nonunitary but energy–conserving; collapse rather than phase dephasing.
This work Unitary spectral deformation $H^{1+i\beta}$ producing oscillatory suppression Closed system; purely dynamical phases; no stochasticity; relies on spectral functional calculus.
Table 2: Comparison with non–environmental intrinsic decoherence proposals.
8.2 Environmental / open–system models
Standard environment–induced decoherence is modeled by master equations of Lindblad type, Caldeira–Leggett oscillator baths, quantum trajectories, and, more recently, non–Hermitian Hamiltonian deformations.
Lindblad master equations describe dissipative evolution of a subsystem coupled to a bath, yielding exponential damping of off–diagonal density–matrix elements, and constitute the canonical form of Markovian open–system dynamics [8]. Caldeira–Leggett–type models implement an oscillator bath linearly coupled to the system, leading to Brownian motion and friction with reduced dynamics that is nonunitary once the bath is traced out [9]. Quantum–trajectory approaches condition evolution on measurement records, with stochastic jumps and nonlinear state updates, providing an unraveling of Lindblad dynamics [8].
Non–Hermitian Hamiltonian deformations in the sense of Matsoukas–Roubeas and collaborators [10, 11] encode energy diffusion and dephasing via effective non–Hermitian generators $H-i\Gamma$, related to open Markovian dynamics, imperfect timekeeping, and inverse $TT\bar{T}$ deformations. In this framework, the imaginary parts of the eigenvalues directly produce decay, and one often restores norm via nonlinear renormalization of the state (BNGL evolution) [10]. In related work, unitarity breaking in self–averaging spectral form factors is analyzed using mixed–unitary channels and non–Hermitian Hamiltonians [11].
All of these models either explicitly break unitarity (reduced dynamics) or require nonlinear renormalization of the state. They differ from the present proposal, which preserves exact unitarity at the level of the full evolution operator.
Model Mechanism Comments
Lindblad master equations [8] Dissipative terms; environment–induced decoherence Exponential damping; completely positive trace–preserving maps.
Caldeira–Leggett [9] Coupling to oscillator bath Reduced dynamics nonunitary; bath traced out.
Quantum trajectories [8] Conditioning on measurement outcomes Stochastic jumps; nonlinear normalization.
Matsoukas–Roubeas / del Campo non–Hermitian deformations [10, 11] Non–Hermitian $H-i\gamma H^{2}$; energy diffusion; inverse $TT\bar{T}$ Open–system origin; BNGL normalization; imaginary eigenvalues give decay.
This work No environment; no dissipation; unitary evolution under $H^{1+i\beta}$ Decoherence arises from non–stationary–phase suppression of spectral interference.
Table 3: Comparison with environmental decoherence models and non–Hermitian deformations.
8.3 Novelty table: this work vs. existing frameworks
The following table highlights key structural differences between the present mechanism and existing approaches:
Feature Existing models This work
Unitarity of evolution Typically broken (reduced dynamics) or stochastically modified Exactly unitary
Source of decoherence Noise, dissipation, collapse, coarse–graining Oscillatory spectral phase $E^{i\beta}$
Mathematical structure Master equations, stochastic terms, $E\mapsto f(E)$ Spectral exponentiation $H^{1+i\beta}$
Environment required Usually yes No
Form of decoherence Exponential damping, collapse, diffusion Non–stationary–phase decay $\sim 1/|\beta|$
Modification of Born rule Sometimes (collapse models) Never
Novel parameter Collapse rates, bath couplings Single dimensionless $\beta$
Table 4: Structural novelty of the imaginary–order spectral deformation mechanism.
8.4 Novelty statement
To the best of our knowledge, no prior model of intrinsic decoherence suppresses interference within strictly unitary dynamics by deforming the generator to an imaginary–order spectral power $H^{1+i\beta}$. Existing intrinsic proposals (Milburn, Diósi–Penrose, GRW/CSL, clock–induced decoherence, Snoke–type energy–conserving collapse) introduce nonunitarity through stochastic or Lindblad–type mechanisms, while environmental models (Lindblad, Caldeira–Leggett, non–Hermitian deformations in the sense of Matsoukas–Roubeas and del Campo) require coupling to external degrees of freedom or non–Hermitian generators with imaginary parts in their spectra. None modifies the Hamiltonian by complex spectral powers or produces decoherence through non–stationary–phase suppression of the oscillatory factor $E^{i\beta}$. This establishes the present mechanism as a new class of unitary yet decohering dynamics rooted entirely in spectral geometry. In this sense, the imaginary–order spectral deformation $H^{1+i\beta}$ defines a new class of intrinsically decohering yet globally unitary dynamics, distinct from both collapse models and standard open–system master equations.
9 Discussion and Outlook
Imaginary–order spectral deformation modifies the dynamical phases while preserving the kinematical structure of quantum mechanics. The Born rule and the Hilbert–space inner product are unaffected, and probabilities assigned to projection operators remain exactly as in the standard theory. The mechanism produces spontaneous decoherence without invoking stochastic collapse, environmental coupling, or real–order fractional dynamics.
The non–stationary–phase estimate shows that interference terms in amplitudes and decoherence functionals decay at least as $\mathcal{O}(1/|\beta|)$ when $|\beta|$ is large. This suggests that $\beta$ is, in principle, experimentally accessible. Precision coherence measurements in trapped ions, superconducting qubits, cold atoms, or optomechanical systems, where environmental decoherence can be carefully modeled and reduced, could constrain the magnitude of $\beta$ by comparing observed coherence decay with expectations from environmental models alone. In this respect, the phenomenology parallels that of collapse models, but the underlying mechanism is entirely different [6, 14].
A conceptual limitation of the present treatment is that $\beta$ has been introduced phenomenologically rather than derived from a microscopic model. It remains an open question whether specific quantum–gravity scenarios predict a definite scale or functional dependence for $\beta$, or whether different frameworks would generate distinct signatures. A systematic classification of admissible spectral deformations compatible with diffeomorphism invariance and effective field–theory principles constitutes a natural next step.
Further work may develop a more detailed phenomenology, including explicit models for $\beta$ in terms of quantum–gravity parameters, and investigate the interplay between this spontaneous decoherence and standard environment–induced decoherence. It would also be natural to study continuous–spectrum systems, scattering problems, and quantum chaotic Hamiltonians under the same deformation to explore whether the mechanism has observable consequences beyond simple bound–state examples.
Acknowledgments
The author thanks colleagues in quantum foundations and quantum gravity whose discussions helped refine the presentation of imaginary–order spectral deformations and their relation to existing decoherence models.